OpenStudy (anonymous):

Calculate the maximum speed of a emitted electron:Where the Maximum Kinetic energy of the emitted electron is 2.28*10,^-19J

OpenStudy (anonymous):

Ek= 0.5mv^2 can you figure it out from dere?

OpenStudy (anonymous):

\[K.E.= \frac{1}{2}mV^2\] where \[m=mass of electron=9.1\times 10^{-31}Kg\]

OpenStudy (anonymous):

2.28*10^-19=0.5*9.1*10^-31*v^2 v^2=2*2.28*10^-19/9.1*10^-31=5.0*10^11

OpenStudy (anonymous):

In principle you have to be careful, because electrons with even a very small kinetic energy can be traveling near the speed of light, and hence you would need to use relativistic kinematics. However, when you work this one out, it's only going at about 0.2% of c, so the nonrelativistic formula is OK.

OpenStudy (anonymous):

@Carl_Pham's point of caution should be noted as well

OpenStudy (anonymous):

How we know that a relativistic calculation is needed or not for a particle? ANS: Whenever "pc" (momentum*speed of light) is much smaller than the rest mass of that particle then we don't need relativistic calculation, otherwise we should use relativistic calculation.